Suppose the market demand for playing cards is given by the equation

Q = 6,000,000 – 1,000,000P where Q is the number of decks of cards demanded each year and P is the price in dollars. For a price increase from $2 to $3 per deck, what is the arc price elasticity?


1
Expert's answer
2021-05-03T10:55:27-0400

Solution:

Arc price elasticity of demand ="\\frac{\\%\\;change\\; in\\; quantity\\; demanded}{\\%\\; change\\; in\\; price}\n\u200b"


Derive quantity demanded:

Quantity demanded at $2 = 6,000,000 – 1,000,000(2) = 6,000,000 – 2,000,000 = 4,000,000

Quantity demanded at $3 = 6,000,000 – 1,000,000(3) = 6,000,000 – 3,000,000 = 3,000,000

 

% change in qty demanded = "\frac{Q_{2} -Q_{1}}{(Q_{1}+Q_{2})/2 } \times 100"


"=\frac{3000000 -4000000}{(4000000+300000)/2 } \times 100 = \frac{-1000000}{3500000} \times 100 = -28.57\%"


% change in price ="=\frac{3 -2}{(2+3)/2 } \times 100 = \frac{1}{2.5} \times 100 = 40\%"


Arc price elasticity of demand ="=\frac{-28.57\%}{40\%} = -0.71"

Arc price elasticity of demand = 0.71

Therefore, it is price inelastic since it is below 1. That is a change in price causes a smaller percentage change in demand.